# Real-time Bayesian inference at extreme scale:

### v1

@@ I Justification for ACM Gordon Bell Prize
Fastest time-to-solution of a PDE-based Bayesian inverse problem with 1 billion parameters in 0.2 seconds, a ten-billion-fold speedup over SoA. Largest-to-date unstructured mesh FE simulation with 55.5 trillion DOF on 43,520 GPUs, with 92% weak and 79% strong parallel efficiencies in scaling over a 128$\times$ increase of GPUs on the full-scale El Capitan system.
@@ III-A Tsunami forecasting and the Cascadia subduction zone
Figure 1: (a) Topobathymetric map of the Cascadia subduction zone (CSZ) based on GEBCO’s gridded bathymetric data set. The shaded area illustrates the potentially “locked” portion of the megathrust interface. The red line marks the trench where the subducting Explorer, Juan de Fuca, and Gorda plates begin their descent beneath the North American Plate. (b)–(c) Snapshots of vertical seafloor uplift and seafloor velocity from a physics-based 3D dynamic rupture and seismic wave propagation computation of a magnitude 8.7 earthquake scenario spanning the full margin of the CSZ. (d) Representative block of a 3D multi-block hexahedral mesh of the CSZ, depicting bathymetry-adapted meshing.
@@
Tsunamis are rare events but can be extremely deadly and cause catastrophic socioeconomic losses. In the past century tsunamis have claimed more than 400,000 lives, and individual tsunamis can result in hundreds of billions of dollars in losses, making them among the costliest natural disasters. State-of-the-art (SoA) tsunami early warning systems [Kamigaichi2009] rely on rapid characterization of source parameters primarily from seismic data [Hirshorn2021], issuing alerts within minutes but based on simplified assumptions about the event. These systems typically infer earthquake moment magnitude and hypocenter location, and then trigger tsunami forecasts using precomputed scenarios or simplified fault models, assuming instantaneous, uniform slip on a predefined fault geometry. This approach is inadequate in the near field [leveque2018cascadia], where destructive tsunami waves can arrive on-shore in under ten minutes. In addition, current models fail to capture the complexities of earthquake rupture dynamics [uphoff2017sumatra], including variable slip distributions and slow rupture speeds. This can lead to delayed, missed, or false warnings, especially when early seismic data misrepresent the true tsunami potential, such as during the 2011 Tōhoku, Japan tsunami or the 2024 Cape Mendocino, California earthquake, which did not cause a tsunami, despite five million people receiving evacuation alerts. To improve warning capabilities, real-time ocean bottom pressure data and GPS measurements can be integrated directly to constrain tsunami sources, without relying on assumed fault geometry or magnitude [Tsushima2009, Crowell2024].
@@
SoA tsunami simulations typically solve the nonlinear shallow water equations, sometimes with added dispersive terms or empirical corrections, to model wave generation, propagation, and run-up. These models are widely used in tsunami hazard assessment and early warning but are known to omit important physical processes such as wave dispersion, 3D hydrodynamic interactions with bathymetry and coastal infrastructure, and the dynamics of the coupled Earth–ocean system [abrahams2023comparison]. Large-scale tsunami simulations using high-resolution bathymetry are constrained by simplifications in source modeling and may require minutes to hours per run, limiting their use in real-time settings [Behrens2021].
@@
We present a full-physics Bayesian inversion framework—a so-called digital twin [DigitalTwinsNASEM]—that enables real-time, data-driven tsunami forecasting with dynamic adaptivity to complex source behavior. Our inversion framework employs ocean bottom pressure data, along with the 3D coupled acoustic–gravity wave equations, to infer earthquake-induced spatiotemporal seafloor motion in real time. The solution of this Bayesian inverse problem then provides the boundary forcing to forward propagate the tsunamis toward coastlines and issue forecasts of wave heights at specified locations, with quantified uncertainties. We apply our framework to the Cascadia subduction zone (CSZ), which stretches 1000 km from northern California to British Columbia (Fig. 1). The CSZ has been eerily quiet for over 300 years but is considered overdue for a magnitude 8.0–9.0 megathrust earthquake, with paleoseismic and geodetic evidence indicating a recurrence interval of roughly 250 to 500 years [Atwater1995, Goldfinger2012]. In addition to recently proposed future offshore deployments (e.g., by SZ4D), the NEPTUNE cabled observatory [Barnes2015] has provided continuous offshore observations since 2009 across the northern Cascadia margin, recording several tsunamis and offering valuable data to inform optimal sensor placement.
@@ III-B Bayesian inverse problems
Our work addresses a critically important class of problems that has received little attention in the Gordon Bell Prize competition: inverse problems [Vogel02, ghattas2021learning], in particular in the Bayesian framework [stuart2010inverse, KaipioSomersalo05]. In a forward problem, the input parameters (e.g., initial or boundary conditions, coefficients, source terms) are specified; solution of the governing equations, here assumed to be partial differential equations (PDEs), then yields the state variables describing the behavior of the system. However, forward models typically contain uncertain parameters. Often these uncertain parameters are infinite-dimensional fields, i.e., they vary in space (and possibly time). In inverse problems, we seek to infer uncertain parameter fields from measurements or observations of system states.
Infinite-dimensional inverse problems are usually ill-posed: many parameter fields can be found that are consistent with the observational data to within the noise. The classical output least-squares approach to the inverse problem posits a regularization functional that penalizes unwanted features of the parameter field (e.g., magnitude or roughness), yielding a unique solution [Vogel02]. In contrast, the Bayesian approach seeks to characterize the probability that any parameter field is consistent with the data and prior knowledge. The parameter is treated as a random field, and the solution of the inverse problem becomes a probability density, the posterior distribution [KaipioSomersalo05]. As such the Bayesian approach is very powerful.
This power comes at a price, though: computing the Bayesian solution is generally intractable for infinite-dimensional inverse problems governed by PDEs. For example, computing just the mean of the posterior formally requires numerical integration in the discretized parameter dimension (in our case, one billion); each evaluation of the integrand requires solution of the forward problem, which can take hours or more, and numerous forward solutions are required. As a result, various approximations of the posterior, such as the Laplace approximation [ghattas2021learning], must be invoked, and even then massive-scale computation is necessary, e.g. [isaac2015scalable, Bui-ThanhGhattasMartinEtAl13, CuiLeiLiuEtAl2025]. The specific Bayesian inverse problem we address here is the real-time inference and prediction of tsunamis for early warning, which adds a real-time requirement to the challenges described above. Our real-time inversion and forecasting goal would seem futile, since SoA inversion methods (see §IV) would require 50 years on 512 NVIDIA A100 GPUs to solve this problem, whereas we have less than a minute to provide early warning. As we shall see in §V, we are able to achieve real-time performance through a combination of novel parallel inversion algorithms, advanced PDE discretization libraries, and exploitation of leading-edge GPU supercomputers.
@@ III-C Inference of seafloor motion with acoustic–gravity model
In this section, we describe our target inverse problem. The forward tsunami dynamics are modeled by the acoustic–gravity wave equations that couple the propagation of ocean acoustic waves with surface gravity waves. This PDE model is derived by linearization of the conservation of mass and momentum around hydrostatic pressure in the compressible ocean, and the coupling to the surface gravity wave comes from a modified free surface boundary condition at the sea surface [lotto2015tsunami]. The model takes the form of a coupled first-order system in the velocity vector field $\vec{u}(\vec{x},t)$, the pressure field $p(\vec{x},t)$, and the surface gravity wave height $\eta(\vec{x},t)$,
$\rho\,\partial_{t}\vec{u}+\nabla p&=0,&\Omega\times(0,T),\\ K^{-1}\partial_{t}p+\nabla\cdot\vec{u}&=0,&\Omega\times(0,T),\\ p&=\rho g\eta,&\partial\Omega_{\text{s}}\times(0,T),\\ \partial_{t}\eta&=\vec{u}\cdot\vec{n},&\partial\Omega_{\text{s}}\times(0,T),\\ \vec{u}\cdot\vec{n}&=-\partial_{t}b,&\partial\Omega_{\text{b}}\times(0,T),\\ \vec{u}\cdot\vec{n}&=Z^{-1}p,&\partial\Omega_{\text{a}}\times(0,T),\\ \end{aligned}\right.$ (1)
@@
with homogeneous initial conditions. Here, $K$ and $\rho$ are the bulk modulus and density of seawater, $Z=\rho c$ is the acoustic wave impedance, $c=\sqrt{K/\rho}$ is the speed of sound in seawater, $g$ is gravitational acceleration, and $\partial_{t}b$ is the seafloor normal velocity with $b$ the normal displacement. The spatial domain is denoted by $\Omega$ with boundaries $\partial\Omega_{\text{s}}$ (sea surface), $\partial\Omega_{\text{b}}$ (sea bottom), and $\partial\Omega_{\text{a}}$ (lateral, absorbing boundaries); $\vec{n}$ is the outward unit normal; and the temporal domain is $(0,T)$.
@@
The uncertain parameter field $m(\vec{x},t)$, to be inferred from the data, is the spatiotemporal seafloor normal velocity,
$m(\vec{x},t)\coloneq-\partial b(\vec{x},t)/\partial t,\quad(\vec{x},t)\in\partial\Omega_{\text{b}}\times(0,T).$
@@
The observables $d$ are model predictions of data $d^{\text{obs}}$ from pressure sensors mounted on the seafloor, i.e. $d=p(\vec{x}^{\hskip 1.0ptd}_{j},t^{d}_{i})$, where $\vec{x}^{\hskip 1.0ptd}_{j}\in\partial\Omega_{\text{b}}$, $j=1,\ldots,N_{d}$, are the seafloor sensor locations and $t^{d}_{i}\in(0,T)$, $i=1,\ldots,N_{t}^{d}$, are the time instants at which observations are made. The parameter-to-observable (p2o) map $\mathcal{F}$ takes the seafloor velocity $m$ as input, solves the acoustic–gravity wave equations (1), and extracts $d$, the pressures at the sensor locations and observation times. The quantities of interest (QoI) are the surface wave height predictions at locations and times of interest, i.e. $q=\eta(\vec{x}^{\hskip 1.0ptq}_{j},t^{q}_{i})$ where $\vec{x}^{\hskip 1.0ptq}_{j}\in\partial\Omega_{\text{s}}$, $j=1,\ldots,N_{q}$, and $t^{q}_{i}\in(0,T)$, $i=1,\ldots,N_{t}^{q}$, are specific locations and time instants pertinent to tsunami early warning. The parameter-to-QoI (p2q) map $\mathcal{F}_{q}$ takes the inferred seafloor velocity $m$ and forward predicts the QoI $q$.
@@
The Bayesian inverse problem can then be stated as follows: given pressure recordings $d^{\text{obs}}$ from sensors on the seafloor, infer the spatiotemporal seafloor velocity $m$ in the subduction zone and its uncertainty. The tsunami posterior prediction problem is: given inferred parameters $m$ and their uncertainty, forward predict the surface gravity wave heights $q$ and their uncertainty. In §IV, we discuss why this problem is intractable with the current SoA.
@@ IV Current State of the Art
where $\pi_{\text{prior}}(\mathbf{m})$ is the prior probability density of the seafloor velocity parameters $\mathbf{m}$, $\pi_{\text{like}}(\mathbf{d}^{\text{obs}}|\mathbf{m})$ is the likelihood of the observed seafloor pressure data $\mathbf{d}^{\text{obs}}$, given $\mathbf{m}$, and $\pi_{\text{post}}(\mathbf{m}|\mathbf{d}^{\text{obs}})$ is the posterior density reflecting the probability of the parameters conditioned on the data.
Taking a Gaussian prior with mean $\mathbf{m}_{\text{prior}}$ and covariance $\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}$ (here block diagonal, with each block the inverse of an elliptic PDE operator in space representing a Matérn covariance [stuart2010inverse]), along with centered Gaussian additive noise with noise covariance $\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}$, the posterior is then given by:
$\pi_{\text{post}}(\mathbf{m}|\mathbf{d}^{\text{obs}})\!\propto\!\exp\!\left\{\!-\tfrac{1}{2}\|\mathbf{F}\mathbf{m}\!-\!\mathbf{d}^{\text{obs}}\|_{\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}}^{2}\!\!-\!\tfrac{1}{2}\|\mathbf{m}\!-\!\mathbf{m}_{\text{prior}}\|_{\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{-1}}^{2}\!\right\}\!.$
@@
$\mathbf{m}_{\text{map}}\!\coloneq\!\operatorname*{arg\,min}_{\mathbf{m}}\!\tfrac{1}{2}\|\mathbf{F}\mathbf{m}-\mathbf{d}^{\text{obs}}\|_{\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}}^{2}+\tfrac{1}{2}\|\mathbf{m}-\mathbf{m}_{\text{prior}}\|_{\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{-1}}^{2}.$
@@
$\left(\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{-1}\right)\mathbf{m}_{\text{map}}=\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}\mathbf{d}^{\text{obs}}+\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{-1}\mathbf{m}_{\text{prior}},$ (2)
@@
where $\mathbf{H}\coloneq(\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{-1})$ is the Hessian of the negative log-posterior and $\mathbf{F}^{*}$ is the adjoint of the p2o map.
@@
The posterior is then a Gaussian centered at the MAP point, i.e. $\pi_{\text{post}}(\mathbf{m}|\mathbf{d}^{\text{obs}})\propto\mathcal{N}(\mathbf{m}_{\text{map}},\mathbf{\Gamma}_{\hskip-2.0pt\text{post}})$, with posterior covariance
@@
$\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}\coloneq\mathbf{H}^{-1}=\left(\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{-1}\right)^{-1}.$ (3)
@@
Furthermore, the posterior predictive of the wave height QoI $\mathbf{q}$, which depend linearly on the parameters via the discrete p2q map $\mathbf{F}_{\hskip-1.0ptq}$, can be determined as the Gaussian probability density $\pi_{\text{post}}(\mathbf{q}|\mathbf{d}^{\text{obs}})\propto\mathcal{N}\left(\mathbf{q}_{\text{map}},\mathbf{\Gamma}_{\hskip-2.0pt\text{post}(q)}\right)$, where $\mathbf{q}_{\text{map}}=\mathbf{F}_{\hskip-1.0ptq}\mathbf{m}_{\text{map}}$ and $\mathbf{\Gamma}_{\hskip-2.0pt\text{post}(q)}=\mathbf{F}_{\hskip-1.0ptq}\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}\mathbf{F}_{\hskip-1.0ptq}^{*}$.
@@
Writing expressions (2) for the posterior mean $\mathbf{m}_{\text{map}}$ and (3) for the posterior covariance $\mathbf{\Gamma}_{\hskip-2.0pt\text{post}}$ is easy; computing them is a monumental challenge. The coefficient matrix of (2), the Hessian $\mathbf{H}$, is a dense matrix of dimension of the parameters, here $10^{9}\times 10^{9}$. Constructing it efficiently via $\mathbf{F}$ requires as many adjoint wave propagation solutions—here, each requiring $\sim$1 hour on 512 NVIDIA A100 GPUs—as there are spatiotemporal data, here $\sim$250,000; this amounts to 25 years. Even if $\mathbf{H}$ could be constructed and stored (4 exabytes), factoring it would require 10 years on a sustained 1 EFLOP/s machine. SoA methods for large-scale inverse problems employ matrix-free conjugate gradients (CG), preconditioned by the prior covariance (thus involving elliptic PDE solves), and yield convergence in a number of iterations of the order of the number of eigenvalues of the prior-preconditioned Hessian of the negative log likelihood, $\widetilde{\mathbf{H}}_{\text{like}}\coloneq\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{1/2}\mathbf{F}^{*}\mathbf{\Gamma}_{\hskip-2.0pt\text{noise}}^{-1}\mathbf{F}\mathbf{\Gamma}_{\hskip-2.0pt\text{prior}}^{1/2}$, that are above unity [ghattas2021learning]. For highly ill-posed inverse problems, the eigenvalues decay rapidly and $\widetilde{\mathbf{H}}_{\text{like}}$ has low effective rank, and thus CG converges rapidly [ghattas2021learning]. Moreover, a low-rank approximation of $\widetilde{\mathbf{H}}_{\text{like}}$ computed via a matrix-free randomized eigensolver, along with the Sherman–Morrison–Woodbury formula, can be used to compactly represent the posterior covariance, sample from the posterior distribution, and compute the posterior variance field, as has been done for some large-scale Bayesian inverse problems [isaac2015scalable, Bui-ThanhGhattasMartinEtAl13].

### v2

@@ I Justification for ACM Gordon Bell Prize
[…]
@@ III-A Tsunami forecasting and the Cascadia subduction zone
Figure 1: (a) Topobathymetric map of the Cascadia subduction zone (CSZ) based on GEBCO’s (General Bathymetric Chart of the Oceans) 15-arc-second gridded bathymetric data set. The shaded area illustrates the potentially “locked” portion of the megathrust interface. The red line marks the trench where the subducting Explorer, Juan de Fuca, and Gorda plates begin their descent beneath the North American Plate. (b)–(c) Snapshots of vertical seafloor uplift and seafloor velocity from a physics-based 3D dynamic rupture and seismic wave propagation computation of a magnitude 8.7 earthquake scenario spanning the full margin of the CSZ. (d) Representative block of a 3D multi-block hexahedral mesh of the CSZ, depicting bathymetry-adapted meshing.
@@
Tsunamis are rare events but can be extremely deadly and cause catastrophic socioeconomic losses. In the past century tsunamis have claimed more than 400,000 lives, and individual tsunamis can result in hundreds of billions of dollars in damage, making them among the costliest natural disasters. State-of-the-art (SoA) tsunami early warning systems [1] rely on rapid characterization of source parameters primarily from seismic data [2], issuing alerts within minutes but based on simplified assumptions about the event. These systems typically infer earthquake moment magnitude and hypocenter location, and then trigger tsunami forecasts using precomputed scenarios or simplified fault models, assuming instantaneous, uniform slip on a predefined fault geometry. This approach is inadequate in the near field [3], where destructive tsunami waves can arrive on-shore in under ten minutes. In addition, current models fail to capture the complexities of earthquake rupture dynamics [4], including variable slip distributions or slow rupture speeds. These factors can lead to delayed, missed, or false warnings, especially when early seismic data misrepresent the true tsunami potential, such as during the 2011 Tōhoku, Japan tsunami or the 2024 Cape Mendocino, California earthquake, which did not cause a tsunami, despite five million people receiving evacuation alerts. To improve warning capabilities, real-time ocean bottom pressure data and GPS measurements can be integrated directly to constrain tsunami sources, without relying on assumed fault geometry or magnitude [5, 6].
@@
SoA tsunami simulations typically solve the nonlinear shallow water equations, sometimes with added dispersive terms or empirical corrections, to model wave generation, propagation, and run-up. These models are widely used in tsunami hazard assessment and early warning but are known to omit important physical processes such as wave dispersion, 3D hydrodynamic interactions with bathymetry and coastal infrastructure, and the dynamics of the coupled Earth–ocean system [7]. Large-scale tsunami simulations using high-resolution bathymetry are constrained by simplifications in source modeling and may require minutes to hours per run, limiting their use in real-time settings [8].
@@
We present a full-physics Bayesian inversion framework—a so-called digital twin [9]—that enables real-time, data-driven tsunami forecasting with dynamic adaptivity to complex source behavior. Our inversion framework employs ocean bottom pressure data, along with the 3D coupled acoustic–gravity wave equations, to infer earthquake-induced spatiotemporal seafloor motion in real time. The solution of this Bayesian inverse problem then provides the boundary forcing to forward propagate the tsunamis toward coastlines and issue forecasts of wave heights at specified locations with quantified uncertainties. We apply our framework to the Cascadia subduction zone (CSZ), which stretches 1000 km from northern California to British Columbia (Fig. 1). The CSZ has been eerily quiet for over 300 years but is considered overdue for a magnitude 8.0–9.0 megathrust earthquake, with paleoseismic and geodetic evidence indicating a recurrence interval of roughly 250 to 500 years [10, 11]. In addition to recently proposed future offshore deployments (e.g., by SZ4D (Subduction Zones in four Dimensions)), the NEPTUNE cabled observatory [12] has provided continuous offshore observations since 2009 across the northern Cascadia margin, recording several tsunamis and offering valuable data to inform optimal sensor placement.
@@ III-B Bayesian inverse problems
Our work addresses a critically important class of problems that has received little attention in the Gordon Bell Prize competition: inverse problems [13, 14], in particular in the Bayesian framework [15, 16]. In a forward problem, the input parameters (e.g., initial or boundary conditions, coefficients, source terms) are specified; solution of the governing equations, here assumed to be partial differential equations (PDEs), then yields the state variables describing the behavior of the system. However, forward models typically contain uncertain parameters. Often these uncertain parameters are infinite-dimensional fields, i.e., they vary in space (and possibly time). In inverse problems, we seek to infer uncertain parameter fields from measurements or observations of system states.
Infinite-dimensional inverse problems are usually ill-posed: many parameter fields are consistent with the observational data to within the noise. The classical output least-squares approach to the inverse problem posits a regularization functional that penalizes unwanted features of the parameter field (e.g., magnitude or roughness), yielding a unique solution [13]. In contrast, the Bayesian approach seeks to characterize the probability that any parameter field is consistent with the data and prior knowledge. The uncertain parameter is treated as a random field, and the solution of the inverse problem becomes a probability density, the posterior distribution [16]. As such the Bayesian approach is very powerful.
This power comes at a price, though: computing the Bayesian solution is generally intractable for infinite-dimensional inverse problems governed by PDEs. For example, computing just the mean of the posterior formally requires numerical integration in the discretized parameter dimension (in our case, one billion); each evaluation of the integrand requires solution of the forward problem, which can take hours or more, and numerous forward solutions are required. Accordingly, various approximations of the posterior, such as the Laplace approximation [14], must be invoked, and even then massive-scale computation is necessary, e.g. [17, 18, 19]. The specific Bayesian inverse problem we address here is the inference and prediction of tsunamis for early warning, adding a real-time requirement to the challenges described above. Our forecasting goal—to provide early warning in less than a minute—may seem futile, as SoA inversion methods (see §IV) would require 50 years on 512 NVIDIA A100 GPUs to solve this problem. As we shall see in §V, we achieve real-time performance through a combination of novel parallel inversion algorithms, advanced PDE discretization libraries, and exploitation of leading-edge GPU supercomputers.
@@ III-C Inference of seafloor motion with acoustic–gravity model
In this section, we describe our target inverse problem. The forward tsunami dynamics are modeled by the acoustic–gravity wave equations that couple the propagation of ocean acoustic waves with surface gravity waves. This PDE model is derived by linearization of the conservation of mass and momentum around hydrostatic pressure in the compressible ocean, and the coupling to the surface gravity wave comes from a modified free surface boundary condition at the sea surface [20]. The model takes the form of a coupled first-order system in the velocity vector field $\vec{u}(\vec{x},t)$, the pressure field $p(\vec{x},t)$, and the surface gravity wave height $\eta(\vec{x},t)$,
[…]
@@
with homogeneous initial conditions. Here, $K$ and $\rho$ are the bulk modulus and density of seawater, $Z=\rho c$ is the acoustic wave impedance, $c=\sqrt{K/\rho}$ is the speed of sound in seawater, $g$ is gravitational acceleration, and $\partial_{t}b$ is the seafloor inward normal velocity with $b$ the displacement. The spatial domain is denoted by $\Omega$ with boundaries $\partial\Omega_{\text{s}}$ (sea surface), $\partial\Omega_{\text{b}}$ (sea bottom), and $\partial\Omega_{\text{a}}$ (lateral, absorbing boundaries); $\vec{n}$ is the outward unit normal; and the temporal domain is $(0,T)$.
@@
The uncertain parameter field $m(\vec{x},t)$, to be inferred from the data, is the spatiotemporal seafloor inward normal velocity,
$m(\vec{x},t)\coloneq\partial b(\vec{x},t)/\partial t,\quad(\vec{x},t)\in\partial\Omega_{\text{b}}\times(0,T).$
@@
The observables $d$ are model predictions of data $d^{\text{obs}}$ from pressure sensors mounted on the seafloor, i.e. $d=p(\vec{x}^{\thinspace d}_{j},t^{d}_{i})$, where $\vec{x}^{\thinspace d}_{j}\in\partial\Omega_{\text{b}}$, $j=1,\ldots,N_{d}$, are the seafloor sensor locations and $t^{d}_{i}\in(0,T)$, $i=1,\ldots,N_{t}^{d}$, are the time instants at which observations are made. The parameter-to-observable (p2o) map $\mathcal{F}$ takes the seafloor velocity $m$ as input, solves the acoustic–gravity wave equations (1), and extracts $d$, the pressures at the sensor locations and observation times. The quantities of interest (QoI) are the surface wave height predictions at locations and times of interest, i.e. $q=\eta(\vec{x}^{\thinspace q}_{j},t^{q}_{i})$ where $\vec{x}^{\thinspace q}_{j}\in\partial\Omega_{\text{s}}$, $j=1,\ldots,N_{q}$, and $t^{q}_{i}\in(0,T)$, $i=1,\ldots,N_{t}^{q}$, are specific locations and time instants pertinent to tsunami early warning. The parameter-to-QoI (p2q) map $\mathcal{F}_{q}$ takes the inferred seafloor velocity $m$ and forward predicts the QoI $q$.
@@
The Bayesian inverse problem can be stated as follows: given pressure recordings $d^{\text{obs}}$ from sensors on the seafloor, infer the spatiotemporal seafloor velocity $m$ in the subduction zone and its uncertainty. The tsunami posterior prediction problem is: given inferred parameters $m$ and their uncertainty, forward predict the surface gravity wave heights $q$ and their uncertainty. In §IV, we discuss why this problem is intractable with the current SoA.
@@ IV Current State of the Art
where $\pi_{\text{prior}}(\mathbf{m})$ is the prior probability density of the seafloor velocity parameters $\mathbf{m}$, $\pi_{\text{like}}(\mathbf{d}^{\text{obs}}|\mathbf{m})$ is the likelihood of the observed seafloor pressure data $\mathbf{d}^{\text{obs}}$ given $\mathbf{m}$, and $\pi_{\text{post}}(\mathbf{m}|\mathbf{d}^{\text{obs}})$ is the posterior density reflecting the probability of the parameters conditioned on the data.
Taking a Gaussian prior with mean $\mathbf{m}_{\text{prior}}$ and covariance $\mathbf{\Gamma}_{\thinspace\text{prior}}$ (here block diagonal, with each block the inverse of an elliptic PDE operator in space representing a Matérn covariance [15]), along with centered Gaussian additive noise with noise covariance $\mathbf{\Gamma}_{\thinspace\text{noise}}$, the posterior is given by:
$\pi_{\text{post}}(\mathbf{m}|\mathbf{d}^{\text{obs}})\!\propto\!\exp\!\left\{\!-\tfrac{1}{2}\|\mathbf{F}\mathbf{m}\!-\!\mathbf{d}^{\text{obs}}\|_{\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}}^{2}\!\!-\!\tfrac{1}{2}\|\mathbf{m}\!-\!\mathbf{m}_{\text{prior}}\|_{\mathbf{\Gamma}_{\thinspace\text{prior}}^{-1}}^{2}\!\right\}\!.$
@@
$\mathbf{m}_{\text{map}}\!\coloneq\!\operatorname*{arg\,min}_{\mathbf{m}}\!\tfrac{1}{2}\|\mathbf{F}\mathbf{m}-\mathbf{d}^{\text{obs}}\|_{\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}}^{2}+\tfrac{1}{2}\|\mathbf{m}-\mathbf{m}_{\text{prior}}\|_{\mathbf{\Gamma}_{\thinspace\text{prior}}^{-1}}^{2}.$
@@
$\left(\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\thinspace\text{prior}}^{-1}\right)\mathbf{m}_{\text{map}}=\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}\mathbf{d}^{\text{obs}}+\mathbf{\Gamma}_{\thinspace\text{prior}}^{-1}\mathbf{m}_{\text{prior}},$ (2)
@@
where $\mathbf{H}\coloneq(\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\thinspace\text{prior}}^{-1})$ is the Hessian of the negative log-posterior and $\mathbf{F}^{*}$ is the adjoint of the p2o map.
@@
The posterior is then a Gaussian centered at the MAP point, i.e. $\pi_{\text{post}}(\mathbf{m}|\mathbf{d}^{\text{obs}})\propto\mathcal{N}(\mathbf{m}_{\text{map}},\mathbf{\Gamma}_{\thinspace\text{post}})$, with posterior covariance
@@
$\mathbf{\Gamma}_{\thinspace\text{post}}\coloneq\mathbf{H}^{-1}=\left(\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}\mathbf{F}+\mathbf{\Gamma}_{\thinspace\text{prior}}^{-1}\right)^{-1}.$ (3)
@@
Furthermore, the posterior predictive of the wave height QoI $\mathbf{q}$, which depend linearly on the parameters via the discrete p2q map $\mathbf{F}_{\thinspace q}$, can be determined as the Gaussian probability density $\pi_{\text{post}}(\mathbf{q}|\mathbf{d}^{\text{obs}})\propto\mathcal{N}\left(\mathbf{q}_{\text{map}},\mathbf{\Gamma}_{\thinspace\text{post}(q)}\right)$, where $\mathbf{q}_{\text{map}}=\mathbf{F}_{\thinspace q}\mathbf{m}_{\text{map}}$ and $\mathbf{\Gamma}_{\thinspace\text{post}(q)}=\mathbf{F}_{\thinspace q}\mathbf{\Gamma}_{\thinspace\text{post}}\mathbf{F}_{\thinspace q}^{*}$.
@@
Writing expressions (2) for the posterior mean $\mathbf{m}_{\text{map}}$ and (3) for the posterior covariance $\mathbf{\Gamma}_{\thinspace\text{post}}$ is easy; computing them is a monumental challenge. The coefficient matrix of (2), the Hessian $\mathbf{H}$, is a dense matrix of dimension of the parameters, here $10^{9}\times 10^{9}$. Constructing it efficiently via $\mathbf{F}$ requires as many adjoint wave propagation solutions—here, each requiring $\sim$1 hour on 512 NVIDIA A100 GPUs—as there are spatiotemporal data, here $\sim$250,000; this amounts to 25 years. Even if $\mathbf{H}$ could be constructed and stored (4 exabytes), factoring it would require 10 years on a sustained 1 EFLOP/s machine. SoA methods for large-scale inverse problems employ matrix-free conjugate gradients (CG), preconditioned by the prior covariance (thus involving elliptic PDE solves), and yield convergence in a number of iterations of the order of the number of eigenvalues of the prior-preconditioned Hessian of the negative log likelihood, $\widetilde{\mathbf{H}}_{\text{like}}\coloneq\mathbf{\Gamma}_{\thinspace\text{prior}}^{1/2}\mathbf{F}^{*}\mathbf{\Gamma}_{\thinspace\text{noise}}^{-1}\mathbf{F}\mathbf{\Gamma}_{\thinspace\text{prior}}^{1/2}$, that are above unity [14]. For highly ill-posed inverse problems, the eigenvalues decay rapidly and $\widetilde{\mathbf{H}}_{\text{like}}$ has low effective rank, and thus CG converges rapidly [14]. Moreover, a low-rank approximation of $\widetilde{\mathbf{H}}_{\text{like}}$ computed via a matrix-free randomized eigensolver, along with the Sherman–Morrison–Woodbury formula, can be used to compactly represent the posterior covariance, sample from the posterior distribution, and compute the posterior variance field, as has been done for some large-scale Bayesian inverse problems [17, 18].